OpenStudy (klepez):

symplify 6 over the square root of 3 +2

OpenStudy (klepez):

here are the choices \[\frac{ 4 }{ \sqrt{3} }\]

OpenStudy (klepez):

\[12-6\sqrt{3}\]

OpenStudy (klepez):

\[\sqrt{3}-2\]

OpenStudy (klepez):

and 3

OpenStudy (klepez):

this is what the problem looks like \[\frac{ 6 }{ \sqrt{3} +2}\]

OpenStudy (klepez):

@Luis_Rivera ?

OpenStudy (anonymous):

i think there's a full moon out cause I believe he made a mistake with his denominator..

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[\frac{ 6 }{ \sqrt{3}+2 }*\frac{ \sqrt{3}-2 }{ \sqrt{3}-2 }\]\[\frac{ 6(\sqrt{3}-2) }{ 3-4 }=\frac{ 6(\sqrt{3}-2) }{ -1 }= -6(\sqrt{3}-2) \]

OpenStudy (anonymous):

therefore 12 - 6 sqr root 3

OpenStudy (anonymous):

which is B